Lars:

The interior angles formed by the sides of a quadrilateral have measures that sum to 360°. What is the value of x?

Lars:

Shadow:

If all the interior angles of a triangle add up to 180, and the three angles of a triangle are a, b, and c, I can write: a + b + c = 180

Shadow:

Try to apply that logic to this quadrilateral. If all the angles add up to something, how can you then represent that situation?

Shadow:



Lars:

@shadow wrote:
Created with Raphaëlabca + b + c = 180Reply Using Drawing
I still don't get it-

XLynx:

cuz ur low IQ

XLynx:

you have a pigeon brain xd

XLynx:

(3x-6) + (2x) + 88 + 108 = 360

XLynx:

5x + 82 + 108 = 360

XLynx:

5x + 190 = 360

XLynx:

5x = 170

XLynx:

x = 34

XLynx:

dummy

Lars:

@xlynx wrote:
(3x-6) + (2x) + 88 + 108 = 360
Thank you for explaining the math process.

XLynx:

@lars wrote:
@xlynx wrote:
(3x-6) + (2x) + 88 + 108 = 360
Thank you for explaining the math process.
np i just have big brain